# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 3.9 Proofs
see for instance, [CGGH83]. In other words, $v\in{{\bf T}}_{t}L$ is characterised by the fact that the infinitesimal monodromy of $\bar{t}$ along the vector $v$ is still in ${\sf T}$. If the Hodge cycle is general then by 3.9 its Hodge locus is a general leaf scheme and by 3.6, the map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ is surjective. This together with 3.1 finishes the proof. ∎
@@ 3.10 Taylor series of periods over Hodge cycles
(44) $f_{t}:=-x_{0}^{d}+x_{1}^{d}-x_{2}^{d}+x_{3}^{d}+\cdots-x_{n}^{d}+x_{n+1}^{d}-\sum_{\alpha}t_{\alpha}x^{\alpha}=0,$
@@
where $\alpha$ runs through a finite subset $I$ of $\mathbb{N}_{0}^{n+2}$ with $\sum_{i=0}^{n+1}\alpha_{i}=d$. For a rational number $r$ let $[r]$ be the integer part of $r$, that is $[r]\leq r<[r]+1$, and $\{r\}:=r-[r]$. Let also $(x)_{y}:=x(x+1)(x+2)\cdots(x+y-1),\ (x)_{0}:=1$ be the Pochhammer symbol. We compute the Taylor series of the integration of differential forms over monodromies of the algebraic cycle
@@
inside the Fermat variety $X_{0}$. The following has been proved in [Mov21a, Theorem 18.9]
@@
Let $\delta_{t}\in H_{n}(X_{t},\mathbb{Z}),\ t\in({\sf T},0)$ be the monodromy (parallel transport) of the cycle $\delta_{0}:=[{\mathbb{P}}^{\frac{n}{2}}]\in H_{n}(X_{0},\mathbb{Z})$ along a path which connects 00 to $t$. For a monomial $x^{\beta}=x_{0}^{\beta_{0}}x_{1}^{\beta_{1}}x_{2}^{\beta_{2}}\cdots x_{n+1}^{\beta_{n+1}}$ with $k:=\sum_{i=0}^{n+1}\frac{\beta_{i}+1}{d}\in\mathbb{N}$ we have
(45) $\frac{(-1)^{\frac{n}{2}}\cdot d^{\frac{n}{2}+1}\cdot(k-1)!}{(2\pi\sqrt{-1})^{\frac{n}{2}}}{{\int}}_{\delta_{t}}{\rm Resi}\left(\frac{x^{\beta}\Omega}{f^{k}_{t}}\right)={{{\sum}}}_{a:I{}\to\mathbb{N}_{0}}\frac{(-1)^{E_{\beta+a^{*}}}D_{\beta+a^{*}}}{a!}\cdot t^{a},$
where the sum runs through all #$\#I{}$-tuples $a=(a_{\alpha},\ \ \alpha\in I{})$ of non-negative integers such that for $\check{\beta}:=\beta+a^{*}$ we have
(46) $\left\{\frac{\check{\beta}_{2e}+1}{d}\right\}+\left\{\frac{\check{\beta}_{2e+1}+1}{d}\right\}=1,\ \ \ \forall e=0,1,\cdots,\frac{n}{2},$
@@
$D_{\check{\beta}}$ ::=:= ${\prod}_{i=0}^{n+1}\left(\left\{\frac{\check{\beta}_{i}+1}{d}\right\}\right)_{\left[\frac{\check{\beta}_{i}+1}{d}\right]},\ \ \ \ E_{\check{\beta}}:=\sum_{e=0}^{\frac{n}{2}}\left[\frac{\check{\beta}_{2e}+1}{d}\right].$
Note that for two types of $a$ the coefficient of $t^{a}$ in (45) is zero. First, when $\beta+a^{*}$ does not satisfy (46). Second, when an entry of $\beta+a^{*}$ plus one is divisible by $d$ (this is hidden in the definition of $D_{\beta+a^{*}}$). The coefficients of the Taylor series are in $\mathbb{Q}$. The Taylor series (27) is just obtained from (45) by setting $n=2,\ d=4,\ \beta=0,\ k=1$. Since in this case $H^{2,0}$ is one dimensional and it is generated by $\omega:={\rm Resi}\left(\frac{\Omega}{f_{t}}\right)$, the Hodge locus in $V$ corresponding to the homology class of the line $\mathbb{P}^{1}:x_{0}-x_{1}=x_{2}-x_{3}=0$ is given by the zero locus of $\int_{\delta_{t}}\omega$ whose Taylor series at ${t}=0$ is given in (27).
In the following, we use the computer implementation of (46) and its general format in [Mov21a, Theorem 18.9]. The full family of hypersurfaces has too much parameters, and so, one has to consider lower truncation of the Taylor series. Our main goal is to show that the natural generators of the ideal of a Hodge locus, are not necessarily defined over ${\mathfrak{R}}(\delta_{0})$ and one has to invert infinitely many primes. In particular, in 3.1 the natural generators of the ideal might not prove this conjecture, and that is why, in this conjecture we have claimed that the ideal is defined over the ring ${\mathfrak{R}}$ and not its natural generators. For example, we consider the family
@@
$1,2^{84}\cdot 5\cdot 7\cdot 11\cdot 13:x_{2}^{2}x_{3}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{1}x_{2}x_{3}^{2},2^{84}\cdot 11:x_{0}x_{2}x_{3}^{2},2^{8}5\cdot 7\cdot 11\cdot 13:x_{1}^{2}x_{3}^{2},2^{86}\cdot 11\cdot 13:x_{0}x_{1}x_{3}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{3}^{2},2^{86}\cdot 7\cdot 11\cdot 13:$
$x_{1}x_{2}^{2}x_{3},2^{85}\cdot 7\cdot 11\cdot 13:x_{0}x_{2}^{2}x_{3},2^{85}\cdot 7\cdot 11\cdot 13:x_{1}^{2}x_{2}x_{3},2^{86}\cdot 7\cdot 11\cdot 13:x_{0}x_{1}x_{2}x_{3},2^{85}\cdot 11:x_{0}^{2}x_{2}x_{3},2^{86}\cdot 11:x_{0}x_{1}^{2}x_{3},2^{84}\cdot 11:x_{0}^{2}x_{1}x_{3},2^{85}\cdot 7\cdot 11\cdot 13:$
$x_{1}^{2}x_{2}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{0}x_{1}x_{2}^{2},2^{86}\cdot 11:x_{0}^{2}x_{2}^{2},2^{86}\cdot 11\cdot 13:x_{0}x_{1}^{2}x_{2},2^{85}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{1}x_{2},2^{85}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{1}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{1}^{2}x_{2}^{2}x_{3}^{2},2^{88}\cdot 7\cdot 11\cdot 13:$
@@ 3.11 Hodge-Tate varieties
Hodge conjecture is true for Hodge-Tate varieties $X$, that is, there are algebraic cycles $Z_{i},\ i=1,2,\ldots,a$ in $X$ of dimension $\frac{m}{2}$ such that $[Z_{i}]$’s generate the homology group $H_{m}(X,\mathbb{Q})$.
@@
This is a consequence of the Hodge conjecture, see [DMOS82, Proposition 1.5]. A complex version of the above statement is as follows.
@@
Both Hodge and Grothendieck-Katz conjectures imply 3.7. Hodge conjecture is not known for varieties of Hodge-Tate type and Grothendieck-Katz conjecture for Gauss-Manin connections and its factors is known in [Kat72].
If the Hodge conjecture for Hodge-Tate varieties is true then $\delta_{z}=[Z_{z}]$ and the result follows from a version of [DMOS82, Proposition 1.5] for families of algebraic cycles.
Let $a:=\#\{(p,q)|p+q=m,\ h^{p,q}(X_{t})\not=0\}$. N. Katz in [Kat70, Corollary 7.5, page 383] proves that the $p$-curvature of the Gauss-Manin connection of $X/V$ is nilpotent of order at most $a$. He attributes this theorem to P. Deligne. If $X/V$ is of Hodge-Tate type then $a=1$ and so we know that the $p$-curvature of the Gauss-Manin connection of $X/V$ is zero. The Grothendieck-Katz conjecture for Gauss-Manin connections and its factors is known in [Kat72], and so the solutions to the Gauss-Manin connection as a differential equation are algebraic. These are exactly the periods $\int_{\delta_{z}}\omega$, see 3.3. ∎
As far as the author is aware of it, there is no classification of Hodge-Tate varieties, and 3.5 is as open as the Hodge conjecture itself. Toric varieties are Hodge-Tate varieties, and we can look for complete intersections in Toric varieties which are Hodge-Tate. For instance, let $X$ be a smooth hypersurface of degree $d$ in the weighted projective space $\mathbb{P}^{v_{0},v_{1},\cdots,v_{n+1}}$. By a result of J. Steenbrink in [Ste77] we know that for $v_{0}=1$ if $\frac{n}{2}\leq\frac{\sum_{i=1}^{n+1}v_{i}}{d}$ then $X$ is of Hodge-Tate type. For an example of this, and the resulting algebraic periods see [Mov21a, Section 16.9].
4 Ramanujan vector field
On se propose de donner un dictionnaire heuristique entre énoncés en cohomologie $l$-adique et énoncés en théorie de Hodge. Ce dictionnaire a notamment pour sources […] et la théorie conjecturale des motifs de Grothendieck […]. Jusqu’ici, il a surtout servi a formuler, en théorie de Hodge, des conjectures, et il en a parfois suggéré une démonstrdtion, (P. Deligne in [Del71]).
@@ 4.1 Introduction
(47) ${\sf v}:=(t_{1}^{2}-\frac{1}{12}t_{2})\frac{\partial}{\partial t_{1}}+(4t_{1}t_{2}-6t_{3})\frac{\partial}{\partial t_{2}}+(6t_{1}t_{3}-\frac{1}{3}t_{2}^{2})\frac{\partial}{\partial t_{3}}$
in characteristic $p\not=2,3$. All the algebraic leaves of this vector field over $\mathbb{C}$ are inside the hypersurface $027t_{3}^{2}-t_{2}^{3}=$ 0 and it has a transcendental solution given by the Eisenstein series $a_{1}E_{2},a_{2}E_{4},a_{3}E_{6}$, where $a=(-\frac{1}{12},\frac{1}{12},-\frac{1}{216})$, and that is why it carries this name. We consider it as a vector field in $\mathbb{A}^{3}_{\mathfrak{R}}={\rm Spec}({\mathfrak{R}}[t_{1},t_{2},t_{3}])$, where ${\mathfrak{R}}$ is a ring of characteristic zero with $2$ and $3$ invertible. We usually take ${\mathfrak{R}}=\mathbb{Z}[\frac{1}{6}]$ and so ${\mathfrak{R}}/p{\mathfrak{R}}=\mathbb{F}_{p}$. Let ${\mathfrak{k}}$ be the quotient field of ${\mathfrak{R}}$. In [Mov12] the author gave a moduli space interpretation of the Ramanujan vector field:
@@
${\sf E}$ ::: $zy^{2}-4(x-t_{1}z)^{3}+t_{2}z^{2}(x-t_{1}z)+t_{3}z^{3}=0,$
$[x;y;z]\times(t_{1},t_{2},t_{3})\in\mathbb{P}^{2}\times{\sf T},$
$(\alpha,\omega)$ ::=:= $\left(\left[\frac{dx}{y}\right],\left[\frac{xdx}{y}\right]\right)\ \ \hbox{given in the affine coordinate $z=1$.}$
@@
Moreover, if $\nabla:H^{1}_{\rm dR}({\sf E}/{\sf T})\to\Omega^{1}_{{\sf T}}\otimes_{{\cal O}^{1}_{{\sf T}}}H^{1}_{\rm dR}({\sf E}/{\sf T})$ is the Gauss-Manin connection of ${\sf E}/{\sf T}$ then the Ramanujan vector field is the unique vector field in ${\sf T}$ with the property
@@
Once we master all the preliminaries of 4.1, the proof becomes an easy exercise, see also [Kat73b, Appendix 1, page 158], and we realize that being an elliptic curve does not play a significant role in this theorem. The author took the job of generalizing 4.1 and wrote many papers with the title “Gauss-Manin connection in disguise”. The project has been summarized in the book [Mov22], and it has been so far written over fields of characteristic zero. For arithmetic applications, it seems to be necessary to consider fields of positive characteristic. We redefine ${\sf T}$ to be ${\rm Spec}\mathbb{Z}[t_{1},t_{2},t_{3},\frac{1}{\Delta}]$ which has a similar moduli space interpretation. In this article we prove the following:
@@
It is shown by J.V. Pereira in [Per02], that foliations modulo primes might have algebraic leaves, even though in characteristic zero they do not have such leaves. 4.2 is a manifestation of this phenomena. Most of the ingredients of the proof of 4.2 comes from the articles of Swinnerton-Dyer, Serre and Katz in LNM 350 and they are the building blocks of the theory of $p$-adic modular forms. The integrality of the coefficients of $E_{2},E_{4},E_{6}$ play a main role in the proof of 4.2. For Calabi-Yau varieties such integralities, and in particular the integrality of the mirror map, have been experimentally observed by physicists, and proved for special class of Calabi-Yau varieties, see [Mov17a, Appendix C] and the references therein, but in general it is an open problem. In [Mov17a, AMSY16] we have introduced the theory of Calabi-Yau modular forms and proved a similar statement as in 4.1 for Calabi-Yau threefolds. We strongly believe that 4.2 for Calabi-Yau varieties is equivalent to the $p$-integrality of Calabi-Yau modular forms. In this direction we observe the following. For a moment forget that the Ramanujan vector field has a solution given by the Eisenstein series. Instead, consider ${\sf v}$ as an ordinary differential equation:
(48) ${\rm R}:\left\{\begin{array}[]{l}\dot{t}_{1}=t_{1}^{2}-\frac{1}{12}t_{2}\\ \dot{t}_{2}=4t_{1}t_{2}-6t_{3}\\ \dot{t}_{3}=6t_{1}t_{3}-\frac{1}{3}t_{2}^{2}\end{array}\right..$
We write each $t_{i}$ as a formal power series in $q$, $t_{i}=\sum_{n=0}^{\infty}t_{i,n}q^{n},\ i=1,2,3$ and substitute in the above differential equation with $\dot{t}_{i}=-q\frac{\partial t_{i}}{\partial q}$ and the initial values $t_{1}=-\frac{1}{12}(1-24q+\dots)$. It turns out that all $t_{i,n}$ can be computed recursively, however, by recursion we can at most claim $t_{i,n}\in\mathbb{Q}$, for more details see [Mov12, Section 4.3].
@@
Let $p\not=2,3$ be a prime. There is a curve in ${\sf T}/\mathbb{F}_{p}$ passing through $a:=(-\frac{1}{12},\frac{1}{12},-\frac{1}{216})$ and tangent both to the Ramanujan vector field ${\sf v}$ and $b=(2,20,\frac{7}{3})$ at $a$, if and only if the solution $t_{1},t_{2},t_{3}$ of ${\sf v}$ described above is $p$-integral, that is, $p$ does not appear in the denominator of $t_{i,n}$’s.
The author strongly believes that the existence of algebraic solutions of modular vector fields as in [Mov22] and generalizations of 4.3 can be proved by available methods in algebraic geometry. This will give a purely geometric method for proving the integrality of Calabi-Yau modular forms without computing periods explicitly. Despite the fact that in this article we heavily use the Eisenstein series $E_{p-1},\ E_{p+1}$ for $p$ a prime number, and these objects in the framework of Calabi-Yau varieties do not exists or not yet discovered, we have formulated 4.2, Item 2 which describes the curve in 4.1 using only ${\sf v}$ and primary decomposition of ideals.
We would also like to announce the following theorem which suggests that there might be a heuristic dictionary for properties of foliations over $\mathbb{C}$ and properties of foliations in characteristic $p$. This is mainly inspired by a similar work in Hodge theory introduced by P. Deligne in [Del71].
@@
A similar theorem has been proved for the Ramanujan vector field in $\mathbb{C}^{3}$ in [Mov08, Theorem 1]. It has a real analytic first integral $f$ in $\mathbb{C}^{3}\backslash\{\Delta=0\}$ and ${\sf v}$ restricted to $f=0$ has also a real analytic first integral.
@@
Figure 1: Leaves and first integral

### v2

@@ 3.9 Proofs
see for instance, [CGGH1983]. In other words, $v\in{{\bf T}}_{t}L$ is characterised by the fact that the infinitesimal monodromy of $\bar{t}$ along the vector $v$ is still in $\sf T$}. If the Hodge cycle is general then by 3.9 its Hodge locus is a general leaf scheme and by 3.6, the map $\Theta_{{\sf T},L}\to{{\bf T}}_{t}L$ is surjective. This together with 3.1 finishes the proof. ∎
@@ 3.10 Taylor series of periods over Hodge cycles
(3.18) $f_{t}:=-x_{0}^{d}+x_{1}^{d}-x_{2}^{d}+x_{3}^{d}+\cdots-x_{n}^{d}+x_{n+1}^{d}-\sum_{\alpha}t_{\alpha}x^{\alpha}=0,\ \$
@@
where $\alpha$ runs through a finite subset $I$ of $\mathbb{N}_{0}^{n+2}$ with $\sum_{i=0}^{n+1}\alpha_{i}=d$. For a rational number $r$ let $[r]$ be the integer part of $r$, that is $r]\leq r<[r]+1$, and $\{r\}:=r-[r]$. Let also $(x)_{y}:=x(x+1)(x+2)\cdots(x+y-1),\ (x)_{0}:=1$ be the Pochhammer symbol. We compute the Taylor series of the integration of differential forms over monodromies of the algebraic cycle
@@
inside the Fermat variety $X_{0}$. The following has been proved in [ho13, Theorem 18.9]
@@
Let $\delta_{t}\in H_{n}(X_{t},\mathbb{Z}),\ t\in({\sf T},0)$ be the monodromy (parallel transport) of the cycle $\delta_{0}:=[{\mathbb{P}}^{\frac{n}{2}}]\in H_{n}(X_{0},\mathbb{Z})$ along a path which connects 0 to $t$. For a monomial $x^{\beta}=x_{0}^{\beta_{0}}x_{1}^{\beta_{1}}x_{2}^{\beta_{2}}\cdots x_{n+1}^{\beta_{n+1}}$ with $k:=\sum_{i=0}^{n+1}\frac{\beta_{i}+1}{d}\in\mathbb{N}$ we have
(3.19) $\frac{(-1)^{\frac{n}{2}}\cdot d^{\frac{n}{2}+1}\cdot(k-1)!}{(2\pi\sqrt{-1})^{\frac{n}{2}}}{{\int}}_{\delta_{t}}{\rm Resi}\left(\frac{x^{\beta}\Omega}{f^{k}_{t}}\right)={{{\sum}}}_{a:I{}\to\mathbb{N}_{0}}\frac{(-1)^{E_{\beta+a^{*}}}D_{\beta+a^{*}}}{a!}\cdot t^{a},$
where the sum runs through all #I\#I{}-tuples $a=(a_{\alpha},\ \ \alpha\in I{})$ of non-negative integers such that for $\check{\beta}:=\beta+a^{*}$ we have
(3.20) $\left\{\frac{\check{\beta}_{2e}+1}{d}\right\}+\left\{\frac{\check{\beta}_{2e+1}+1}{d}\right\}=1,\ \ \ \forall e=0,1,\cdots,\frac{n}{2},$
@@
$D_{\check{\beta}}$ $:=$ ${\prod}_{i=0}^{n+1}\left(\left\{\frac{\check{\beta}_{i}+1}{d}\right\}\right)_{\left[\frac{\check{\beta}_{i}+1}{d}\right]},\ \ \ \ E_{\check{\beta}}:=\sum_{e=0}^{\frac{n}{2}}\left[\frac{\check{\beta}_{2e}+1}{d}\right].$
Note that for two types of $a$ the coefficient of $t^{a}$ in (3.19) is zero. First, when $\beta+a^{*}$ does not satisfy (3.20). Second, when an entry of $\beta+a^{*}$ plus one is divisible by $d$ (this is hidden in the definition of $D_{\beta+a^{*}}$). The coefficients of the Taylor series are in $\mathbb{Q}$. The Taylor series (3.1) is just obtained from (3.19) by setting $n=2,\ d=4,\ \beta=0,\ k=1$. Since in this case $H^{2,0}$ is one dimensional and it is generated by $\omega:={\rm Resi}\left(\frac{\Omega}{f_{t}}\right)$, the Hodge locus in $V$ corresponding to the homology class of the line $\mathbb{P}^{1}:x_{0}-x_{1}=x_{2}-x_{3}=0$ is given by the zero locus of $\int_{\delta_{t}}\omega$ whose Taylor series at t=0{t}=0 is given in (3.1).
In the following, we use the computer implementation of (3.20) and its general format in [ho13, Theorem 18.9]. The full family of hypersurfaces has too much parameters, and so, one has to consider lower truncation of the Taylor series. Our main goal is to show that the natural generators of the ideal of a Hodge locus, are not necessarily defined over ${\mathfrak{R}}(\delta_{0})$ and one has to invert infinitely many primes. In particular, in 3.1 the natural generators of the ideal might not prove this conjecture, and that is why, in this conjecture we have claimed that the ideal is defined over the ring $\mathfrak{R}$} and not its natural generators. For example, we consider the family
@@
$1,2^{84}\cdot 5\cdot 7\cdot 11\cdot 13:x_{2}^{2}x_{3}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{1}x_{2}x_{3}^{2},2^{84}\cdot 11:x_{0}x_{2}x_{3}^{2},2^{8}5\cdot 7\cdot 11\cdot 13:x_{1}^{2}x_{3}^{2},2^{86}\cdot 11\cdot 13:$
$x_{0}x_{1}x_{3}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{3}^{2},2^{86}\cdot 7\cdot 11\cdot 13:x_{1}x_{2}^{2}x_{3},2^{85}\cdot 7\cdot 11\cdot 13:x_{0}x_{2}^{2}x_{3},2^{85}\cdot 7\cdot 11\cdot 13:x_{1}^{2}x_{2}x_{3},2^{86}\cdot 7\cdot 11\cdot 13:$
$x_{0}x_{1}x_{2}x_{3},2^{85}\cdot 11:x_{0}^{2}x_{2}x_{3},2^{86}\cdot 11:x_{0}x_{1}^{2}x_{3},2^{84}\cdot 11:x_{0}^{2}x_{1}x_{3},2^{85}\cdot 7\cdot 11\cdot 13:x_{1}^{2}x_{2}^{2},2^{86}\cdot 7\cdot 11\cdot 13:$
$x_{0}x_{1}x_{2}^{2},2^{86}\cdot 11:x_{0}^{2}x_{2}^{2},2^{86}\cdot 11\cdot 13:x_{0}x_{1}^{2}x_{2},2^{85}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{1}x_{2},2^{85}\cdot 7\cdot 11\cdot 13:x_{0}^{2}x_{1}^{2},2^{86}\cdot 7\cdot 11\cdot 13:$
$x_{0}^{2}x_{1}^{2}x_{2}^{2}x_{3}^{2},2^{88}\cdot 7\cdot 11\cdot 13:$
@@ 3.11 Hodge-Tate varieties
Hodge conjecture is true for Hodge-Tate varieties $X$, that is, there are algebraic cycles $Z_{i},\ i=1,2,\ldots,a$ in $X$ of dimension $\frac{m}{2}$ such that $Z_{i}]$’s generate the homology group $H_{m}(X,\mathbb{Q})$.
@@
This is a consequence of the Hodge conjecture, see [dmos, Proposition 1.5]. A complex version of the above statement is as follows.
@@
Both Hodge and Grothendieck-Katz conjectures imply 3.7. Hodge conjecture is not known for varieties of Hodge-Tate type and Grothendieck-Katz conjecture for Gauss-Manin connections and its factors is known in [Katz1972].
If the Hodge conjecture for Hodge-Tate varieties is true then $\delta_{z}=[Z_{z}]$ and the result follows from a version of [dmos, Proposition 1.5] for families of algebraic cycles.
Let $a:=\#\{(p,q)|p+q=m,\ h^{p,q}(X_{t})\not=0\}$. N. Katz in [Katz1970, Corollary 7.5, page 383] proves that the $p$-curvature of the Gauss-Manin connection of $X/V$ is nilpotent of order at most $a$. He attributes this theorem to P. Deligne. If $X/V$ is of Hodge-Tate type then $a=1$ and so we know that the $p$-curvature of the Gauss-Manin connection of $X/V$ is zero. The Grothendieck-Katz conjecture for Gauss-Manin connections and its factors is known in [Katz1972], and so the solutions to the Gauss-Manin connection as a differential equation are algebraic. These are exactly the periods $\int_{\delta_{z}}\omega$, see 3.3. ∎
As far as the author is aware of it, there is no classification of Hodge-Tate varieties, and 3.5 is as open as the Hodge conjecture itself. Toric varieties are Hodge-Tate varieties, and we can look for complete intersections in Toric varieties which are Hodge-Tate. For instance, let $X$ be a smooth hypersurface of degree $d$ in the weighted projective space $\mathbb{P}^{v_{0},v_{1},\cdots,v_{n+1}}$. By a result of J. Steenbrink in [st77] we know that for $v_{0}=1$ if $\frac{n}{2}\leq\frac{\sum_{i=1}^{n+1}v_{i}}{d}$ then $X$ is of Hodge-Tate type. For an example of this, and the resulting algebraic periods see [ho13, Section 16.9].
Chapter 4 Ramanujan vector field
On se propose de donner un dictionnaire heuristique entre énoncés en cohomologie $l$-adique et énoncés en théorie de Hodge. Ce dictionnaire a notamment pour sources […] et la théorie conjecturale des motifs de Grothendieck […]. Jusqu’ici, il a surtout servi a formuler, en théorie de Hodge, des conjectures, et il en a parfois suggéré une démonstrdtion, (P. Deligne in [de71-0]).
@@ 4.1 Introduction
(4.1) ${\sf v}:=(t_{1}^{2}-\frac{1}{12}t_{2})\frac{\partial}{\partial t_{1}}+(4t_{1}t_{2}-6t_{3})\frac{\partial}{\partial t_{2}}+(6t_{1}t_{3}-\frac{1}{3}t_{2}^{2})\frac{\partial}{\partial t_{3}}$
in characteristic $p\not=2,3$. All the algebraic leaves of this vector field over $\mathbb{C}$ are inside the hypersurface $27t_{3}^{2}-t_{2}^{3}=0$ and it has a transcendental solution given by the Eisenstein series $a_{1}E_{2},a_{2}E_{4},a_{3}E_{6}$, where $a=(-\frac{1}{12},\frac{1}{12},-\frac{1}{216})$, and that is why it carries this name. We consider it as a vector field in $\mathbb{A}^{3}_{\mathfrak{R}}={\rm Spec}({\mathfrak{R}}[t_{1},t_{2},t_{3}])$, where $\mathfrak{R}$} is a ring of characteristic zero with $2$ and $3$ invertible. We usually take ${\mathfrak{R}}=\mathbb{Z}[\frac{1}{6}]$ and so ${\mathfrak{R}}/p{\mathfrak{R}}=\mathbb{F}_{p}$. Let $\mathfrak{k}$} be the quotient field of $\mathfrak{R}$}. In [ho14] the author gave a moduli space interpretation of the Ramanujan vector field:
@@
${\sf E}$ $:$ $zy^{2}-4(x-t_{1}z)^{3}+t_{2}z^{2}(x-t_{1}z)+t_{3}z^{3}=0,$
$[x;y;z]\times(t_{1},t_{2},t_{3})\in\mathbb{P}^{2}\times{\sf T},\$
$(\alpha,\omega)$ $:=$ $\left(\left[\frac{dx}{y}\right],\left[\frac{xdx}{y}\right]\right)\ \ \hbox{given in the affine coordinate $z=1$.}$
@@
Moreover, if $\nabla:H^{1}_{\rm dR}({\sf E}/{\sf T})\to\Omega^{1}_{{\sf T}}\otimes_{{\mathcal{O}}^{1}_{{\sf T}}}H^{1}_{\rm dR}({\sf E}/{\sf T})$ is the Gauss-Manin connection of ${\sf E}/{\sf T}$ then the Ramanujan vector field is the unique vector field in $\sf T$} with the property
@@
The moduli space $\sf T$} in [ho2020] is called ”ibiporanga”, a word in Tupi language meaning pretty land. Once we master all the preliminaries of 4.1, the proof becomes an easy exercise, see also [ka73, Appendix 1, page 158], and we realize that being an elliptic curve does not play a significant role in this theorem. The author took the job of generalizing 4.1 and wrote many papers with the title “Gauss-Manin connection in disguise”. The project has been summarized in the book [ho2020], and it has been so far written over fields of characteristic zero. For arithmetic applications, it seems to be necessary to consider fields of positive characteristic. We redefine $\sf T$} to be ${\rm Spec}\mathbb{Z}[t_{1},t_{2},t_{3},\frac{1}{\Delta}]$ which has a similar moduli space interpretation. In this article we prove the following:
@@
It is shown by J.V. Pereira in [Pereira2002], that foliations modulo primes might have algebraic leaves, even though in characteristic zero they do not have such leaves. 4.2 is a manifestation of this phenomena. Most of the ingredients of the proof of 4.2 comes from the articles of Swinnerton-Dyer, Serre and Katz in LNM 350 and they are the building blocks of the theory of $p$-adic modular forms. The integrality of the coefficients of $E_{2},E_{4},E_{6}$ play a main role in the proof of 4.2. For Calabi-Yau varieties such integralities, and in particular the integrality of the mirror map, have been experimentally observed by physicists, and proved for special class of Calabi-Yau varieties, see [GMCD-MQCY3, Appendix C] and the references therein, but in general it is an open problem. In [GMCD-MQCY3, HosseinMurad] we have introduced the theory of Calabi-Yau modular forms and proved a similar statement as in 4.1 for Calabi-Yau threefolds. We strongly believe that 4.2 for Calabi-Yau varieties is equivalent to the $p$-integrality of Calabi-Yau modular forms. In this direction we observe the following. For a moment forget that the Ramanujan vector field has a solution given by the Eisenstein series. Instead, consider $\sf v$} as an ordinary differential equation:
(4.2) ${\rm R}:\left\{\begin{array}[]{l}\dot{t}_{1}=t_{1}^{2}-\frac{1}{12}t_{2}\\ \dot{t}_{2}=4t_{1}t_{2}-6t_{3}\\ \dot{t}_{3}=6t_{1}t_{3}-\frac{1}{3}t_{2}^{2}\end{array}\right.$.
We write each $t_{i}$ as a formal power series in $q$, $t_{i}=\sum_{n=0}^{\infty}t_{i,n}q^{n},\ i=1,2,3$ and substitute in the above differential equation with $\dot{t}_{i}=-q\frac{\partial t_{i}}{\partial q}$ and the initial values $t_{1}=-\frac{1}{12}(1-24q+\dots)$. It turns out that all $t_{i,n}$ can be computed recursively, however, by recursion we can at most claim $t_{i,n}\in\mathbb{Q}$, for more details see [ho14, Section 4.3].
@@
Let $p\not=2,3$ be a prime. The followings are equivalent:
  1. 1.
    There is a curve $C$ in ${\sf T}/\mathbb{F}_{p}$ passing through $a:=(-\frac{1}{12},\frac{1}{12},-\frac{1}{216})$, smooth at $a$, tangent to $b=(2,20,\frac{7}{3})$ at $a$ and tangent to the Ramanujan vector field $\sf v$} and ${\sf v}^{p}={\sf v}$ restricted to $C$.
  2. 2.
    The solution $t_{1},t_{2},t_{3}$ of $\sf v$} described above is $p$-integral, that is, $p$ does not appear in the denominator of $t_{i,n}$’s.
The author strongly believes that the existence of algebraic solutions of modular vector fields as in [ho2020] and generalizations of 4.3 can be proved by available methods in algebraic geometry. This will give a purely geometric method for proving the integrality of Calabi-Yau modular forms without computing periods explicitly. Despite the fact that in this article we heavily use the Eisenstein series $E_{p-1},\ E_{p+1}$ for $p$ a prime number, and these objects in the framework of Calabi-Yau varieties do not exists or not yet discovered, we have formulated 4.2, Item 2 which describes the curve in 4.1 using only $\sf v$} and primary decomposition of ideals.
We would also like to announce the following theorem which suggests that there might be a heuristic dictionary for properties of foliations over $\mathbb{C}$ and properties of foliations in characteristic $p$. This is mainly inspired by a similar work in Hodge theory introduced by P. Deligne in [de71-0].
@@
[…]
@@
Figure 4.1: Leaves and first integral